OpenStudy (anonymous):

list all posible rational roots 3x^3 + 9x- 6 =0

OpenStudy (anonymous):

The cubed term must be 3, so you know the equation must follow: (3x+a)(x+b)(x+c)=0 Now multiply it out for something rather long: \((3x^2+(3b+a)x+ab)(x+c)\)=0 \(3x^3+(3b+a+3c)x^2+[c(3b+a)+ab]x+abc\)=0 The original equation has no \(x^2\), a 9x, and a 6. So these equations must also be true: \(3b+a+3c\) = 0 \([c(3b+a)+ab]\) = 9 \(abc\) = 6 This gets you three equations and three unknowns. Solve for the unknowns and stick it back into the equation.